OpenStudy (anonymous):

Use the intermediate value theorem to determine if h[theta]=1- 3 tan theta has a zero in the interval [0,pi/4]. Use form IVT to answer

OpenStudy (anonymous):

If \(h(\theta)\) is continuous over than interval, you should be able to check to signs on the endpoints, then determine if \(h\) crosses the \(\theta\)-axis.

OpenStudy (anonymous):

So yes, it is continuous, because the domain of the tangent function is \(\left(-\dfrac{\pi}{2},\dfrac{\pi}{2}\right)\). Evaluate the endpoints: \[h(0)=1-3\tan0=1-0=1\\ h\left(\frac{\pi}{4}\right)=1-3\tan\frac{\pi}{4}=1-3=-2\] Since \(h(\theta)\) changes from positive to negative values over the interval, and since it's continuous over this interval, there must be some value \(c\) such that \(h(c)=0\).